# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 3.3. Supersingular modules and the derived adjoints
We must check that that the recollement obtained in 3.13(i) is non-trivial, i.e. that the quotient categories ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$ are non-zero. This will follow by considering supersingular modules. In this subsection, we assume that $k$ is algebraically closed.
@@
Suppose that either $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ (possibly empty product) or that $r_{ss}>0$ and $q>2$. Then there exists a simple supersingular ${\mathcal{H}}$-module of infinite projective dimension.
@@
If $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ then any simple supersingular ${\mathcal{H}}$-module has infinite projective dimension by [19, Theorem 7.7]. So assume that $G$ has type $A_{1}\times\cdots\times A_{1}$ (non-empty product) and that $q>2$. By [19, Theorem 7.7] again, our claim amounts to showing that there exists a $k$-valued character $\xi$ of the group $T(\mathbb{F}_{q})$ such that $S_{\xi}\neq S$ (cf. (2.2)). But since $p$ does not divide the order of $T(\mathbb{F}_{q})$ and as $T(\mathbb{F}_{q})$ is abelian, the intersection of the kernels of all the non-trivial $k$-valued characters of $T(\mathbb{F}_{q})$ is trivial. Let $\alpha$ be a simple root and $s=s_{\alpha}\in S$, and note that the group $T_{s}(\mathbb{F}_{q})$ is non-trivial by our assumption that $q>2$. Then there must exist a character $\xi$ with $\xi|_{T_{s}(\mathbb{F}_{q})}\neq 1$ and so $S_{\xi}\neq S$. ∎
@@
To see this, we use the fact that $T_{\lambda_{0}^{+}}^{n+1}=z^{n}T_{\lambda_{0}^{+}}$ (c.f. [2, Lemma 5.17] and [29, Lemma 2.18]). Now let $f\in{\mathcal{R}^{G}_{M}}({\mathfrak{m}}^{\prime})$. Suppose first that $f\cdot z^{n}=0$. Then for all $h\in{\mathcal{H}}_{M}$, we have
@@
by definition of the map $j_{M}^{+}$ (c.f. (2.4)). Since $T^{M}_{\lambda_{0}^{+}}$ is invertible in ${\mathcal{H}}_{M}$, this means that $f=0$. To get surjectivity, define $g:{\mathcal{H}}_{M}\to{\mathfrak{m}}^{\prime}$ by $g(h)=f(h\cdot(T^{M}_{\lambda_{0}^{+}})^{-n-1})\cdot T_{\lambda_{0}^{+}}$. This is ${\mathcal{H}}_{M^{+}}$-linear because $T^{M}_{\lambda_{0}^{+}}$ is central in ${\mathcal{H}}_{M}$, and we have
@@
[…]
as a square in ${\operatorname{Ho}}({\mathcal{H}})$, we deduce that multiplication by $z^{n}$ induces the zero endomorphism of $Q_{f}{\mathfrak{m}}$ in the homotopy category. Since $Q_{f}{\mathfrak{m}}$ is Gorenstein injective, this means by 2.14 that there is a factorisation
[…]
where $\mathfrak{e}\in{\operatorname{Mod}}({\mathcal{H}})$ is an injective module. By now applying ${\mathcal{R}^{G}_{M}}$ to that commuting triangle, we deduce from the bijectivity of ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})\stackrel{{\scriptstyle\cdot z^{n}}}{{\longrightarrow}}{\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ that ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ is isomorphic to a direct summand of ${\mathcal{R}^{G}_{M}}(\mathfrak{e})$. But since ${\mathcal{R}^{G}_{M}}$ is the right adjoint of an exact functor, the latter module is injective and thus so is ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$. The result now follows immediately because $\mathbf{R}{\mathcal{R}^{G}_{M}}({\mathfrak{m}})\cong{\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ in ${\operatorname{Ho}}({\mathcal{H}}_{M})$. ∎
@@
Similarly to the above proof, take $\lambda_{0}^{\pm}$ as in 2.1 and let $\mathcal{O}^{\pm}=\widetilde{W}\cdot\lambda_{0}^{\pm}$. Then part (i) of the above result holds more generally whenever every element of ${\mathfrak{m}}$ is killed by a power of $z_{\mathcal{O}^{-}}$, and part (ii) holds whenever $z_{\mathcal{O}^{+}}$ acts nilpotently on ${\mathfrak{m}}$. This follows from the above proof and because the abelian analogue of 3.18 holds in that higher level of generality, c.f. the proof of [4, Proposition 4.27].
@@
As another corollary, we deduce that uniformly supersingular modules are Ext-orthogonal to parabolically induced modules in large enough degree without assuming any finiteness condition on these modules:
@@
Let $Q_{f}{\mathfrak{m}}$ be the fibrant replacement of ${\mathfrak{m}}$. We have by the proof of 3.18(ii) that ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ is injective, and so by 3.7(ii)-(iii) we have
@@
Next, we compute the image of all simple modules under the adjoints of parabolic induction at the homotopy level. For that we need the corresponding formula at the level of the module categories, due to Abe:
@@
So we may assume that $P\subseteq P^{\prime}$ and $\Pi(\mathfrak{n})\subseteq\Pi_{Q}\cup\Pi_{P^{\prime}}$. Next, for any standard parabolic $Q\cap M^{\prime}\subseteq Q^{\prime}\subseteq(P\cap M^{\prime})(\mathfrak{n})$ of $M^{\prime}$ with Levi subgroup $L^{\prime}$ we have $e_{{\mathcal{H}}_{w_{0}(L^{\prime})}}(\iota_{0}^{M}\mathfrak{n})\cong\iota_{0}^{L^{\prime}}e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n})$, c.f. [4, Lemma 4.22]. From this and the above we get
@@
where the second, third and fourth isomorphisms hold by 2.3, [29, Lemma 4.9] and (2.5), respectively. By exactness of $\iota_{0}^{w_{0}(M^{\prime})}$ and from the definitions of both $I_{{\mathcal{H}}_{M^{\prime}}}(P\cap M^{\prime},\mathfrak{n},Q\cap M^{\prime})$ and $CI_{{\mathcal{H}}_{M^{\prime}}}(w_{0}(P\cap M^{\prime}),\iota_{0}^{M}\mathfrak{n},w_{0}(Q\cap M^{\prime}))$, it follows that we have an isomorphism
@@
Lemma 3.24.
Let ${\mathfrak{m}}$ be a finite length ${\mathcal{H}}$-module. Then ${\mathfrak{m}}$ is ${\mathcal{R}^{G}_{M}}$-acyclic for any standard Levi subgroup $M$ of $G$.
@@
for all $j\geq 0$, where $({\mathcal{R}^{G}_{M}})^{\prime}$ denotes the right adjoint of parabolic coinduction. By applying 3.7(iii), we deduce that finite length modules are acyclic for $({\mathcal{R}^{G}_{M}})^{\prime}$. Since $({\mathcal{R}^{G}_{M}})^{\prime}\cong\iota_{0}^{w_{0}(M)}\circ\mathcal{R}^{G}_{w_{0}(M)}$, c.f. 2.3(i), and $\iota_{0}^{w_{0}(M)}$ is an exact equivalence, we get that finite length modules are acyclic for $\mathcal{R}^{G}_{w_{0}(M)}$. As $M$ was arbitrary we are done. ∎
@@
One deduces immediately from 3.24, [2, Proposition 5.18] and 3.7(iii) that supersingular modules of finite length are Ext-orthogonal to all parabolically induced modules in all positive degrees, strengthening 3.21 in that case.
@@
Let $M$ be a standard Levi subgroup of $G$ and let ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ be ${\mathcal{R}^{G}_{M}}$-acyclic. Then there is an isomorphism $\mathbf{R}{\mathcal{R}^{G}_{M}}({\mathfrak{m}})\cong{\mathcal{R}^{G}_{M}}(m)$ in ${\operatorname{Ho}}({\mathcal{H}}_{M})$.
@@
where $\mathfrak{e}$ has finite injective dimension. As both ${\mathfrak{m}}$ and $Q_{f}{\mathfrak{m}}$ are ${\mathcal{R}^{G}_{M}}$-acyclic, c.f. 3.7(ii), so is $\mathfrak{e}$. By applying 3.7(iii), we then get isomorphisms
@@
The statement for the left adjoints follows immediately from 3.23. The statement for the right adjoints follows additionally from 3.24 and 3.26. ∎
Using the above, we give a new proof of one direction of [19, Theorem 8.5] without any assumptions on the prime $p$ and without using any Bruhat-Tits theory. This result is not new, in that the proof of the equivalence in loc. cit. only uses their assumption on $p$ in one of the two direction, although they never state this explicitly. Using their restriction on $p$, we can also essentially reduce our aim of studying the isomorphism classes of simple modules in ${\operatorname{Ho}}({\mathcal{H}})$ to the supersingular case.
@@
[…]
@@
(i) Let $L$ denote the Levi subgroup of $Q$. Then we have $0=\mathbf{R}\mathcal{R}^{G}_{L}(I_{\mathcal{H}}(P,\mathfrak{n},Q))\cong I_{{\mathcal{H}}_{L}}(P\cap L,\mathfrak{n},L)$ by 3.27, so that we may assume that $Q=G$. Applying the left adjoint instead, we now have
@@
(ii) First suppose that $Q\neq Q^{\prime}$. Up to swapping them we may assume that $Q\nsubseteq Q^{\prime}$. Write $L^{\prime}$ for the Levi subgroup of $Q^{\prime}$. Then applying $\mathbf{R}\mathcal{R}^{G}_{M^{\prime}}$ to ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$, we obtain $0\cong I_{{\mathcal{H}}_{L^{\prime}}}(P^{\prime}\cap L^{\prime},\mathfrak{n}^{\prime},L^{\prime})$ by 3.27, contradicting our assumption that ${\mathfrak{m}}^{\prime}$ has infinite projective dimension by [19, Theorem 8.5]. Thus $Q=Q^{\prime}$, and applying $\mathbf{R}\mathcal{R}^{G}_{L^{\prime}}$ again we may assume that $Q=Q^{\prime}=G$ as above. Similarly now, if $P\neq P^{\prime}$ then without loss of generality we have $P\subsetneqq P^{\prime}$ and, writing $M^{\prime}$ for the Levi subgroup of $P^{\prime}$, we then obtain a contradiction from 3.27 by applying ${\operatorname{Ho}}(\mathcal{L}^{G}_{M^{\prime}})$. ∎
@@
[…]
@@
First observe that for any standard parabolic subgroup $P^{\prime}=M^{\prime}N^{\prime}$ of $G$, we have that $I_{{\mathcal{H}}_{M^{\prime}}}(P,\mathfrak{n},Q\cap P^{\prime})$ and $I_{{\mathcal{H}}_{M^{\prime}}}(P,\mathfrak{n},Q)$ also have infinite projective dimension, whenever these modules are well-defined. Indeed, this follows from [19, Theorem 8.5] by our assumption on $p$ and because $\mathfrak{s}$ is assumed to have infinite projective dimension. The statement of (ii) then follows from 3.27 and 3.13(iii), and for (i) we get
@@
[…]
@@
[…]
@@
Suppose first that $\mathfrak{s}=I_{\mathcal{H}}(P,\mathfrak{n},Q)$ is a simple subquotient of ${\mathfrak{m}}_{\text{non-ss}}$. If $P(\mathfrak{n})\neq G$ then $\mathfrak{s}$ is in the essential image of $I^{G}_{M(\mathfrak{n})}$. If instead $P(\mathfrak{n})=G$ then $P=B$ since $P\neq G$ and the root system is irreducible, and thus $\mathfrak{s}$ has finite projective dimension by [19, Theorem 8.5] and the fact that ${\mathcal{H}}_{T}$ has finite global dimension (c.f. 3.6). Either way, it follows that $\mathfrak{s}\in{\mathcal{S}}^{G}_{\text{all}}$ and hence that ${\mathfrak{m}}_{\text{non-ss}}\in{\mathcal{S}}^{G}_{\text{all}}$.
@@
When the hypotheses of 3.31 are satisfied we tentatively call the quotient category ${\operatorname{Ho}}({\mathcal{H}})_{\text{all}}$ the supersingular homotopy category of ${\mathcal{H}}$. We will see later that this triangulated category is in fact the homotopy category of a model structure on ${\operatorname{Mod}}({\mathcal{H}})$.
@@
If we write ${\mathcal{T}}$ for ${\operatorname{Ker}}(\mathcal{L}^{G}_{\text{all}})$, then 3.31 tells us that the modules of finite length in ${\mathcal{T}}$ are, up to weak equivalence, the supersingular ones whenever the root system of $G$ is irreducible and $p\nmid\left|\Omega_{\text{tor}}\right|$. Moreover we have ${\mathcal{T}}^{\ddagger}={\mathcal{S}}^{G}_{\text{all}}$ and ${\mathcal{T}}={{}^{\ddagger}}{\mathcal{S}}^{G}_{\text{all}}$. This is similar to the torsion pair constructed at the abelian level in [26] where they showed that when $\mathbb{G}=\text{SL}_{2}$, $p>2$ and $\mathfrak{F}\neq\mathbb{Q}_{p}$, then the supersingular modules are in a torsion pair with a certain class of modules which they call ‘non-supersingular’ and which contains all parabolically induced modules, c.f. Theorems 3.26 & 3.27 in loc. cit. Since our Hom-orthogonality boils down to Ext-orthogonality between suitable replacements at the abelian level, the two questions are not immediately related. But we hope to study elsewhere the relationship between the Ollivier-Schneider torsion pair and the adjoints of parabolic induction.

### v4

@@ 3.3. Supersingular modules and the derived adjoints
We must check that the recollement obtained in 3.13(i) is non-trivial, i.e. that the quotient categories ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$ are non-zero. This will follow by considering supersingular modules. In this subsection, we assume that $k$ is algebraically closed. For the next result, we let
$c:=\min_{s_{\alpha}\in S}|\mu_{\alpha}|$
with $\mu_{\alpha}$ as in (2.4), and note that $c=1$ or $2$.
@@
Suppose that either $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ (possibly empty product) or that $r_{ss}>0$ and $q>2c$. Then there exists a simple supersingular ${\mathcal{H}}$-module of infinite projective dimension.
@@
If $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ then any simple supersingular ${\mathcal{H}}$-module has infinite projective dimension by [21, Theorem 7.7]. So assume that $G$ has type $A_{1}\times\cdots\times A_{1}$ (non-empty product) and that $q>2c$. By [21, Theorem 7.7] again, our claim amounts to showing that there exists a $k$-valued character $\xi$ of the group $T(\mathbb{F}_{q})$ such that $S_{\xi}\neq S$ (cf. (2.2)). But since $p$ does not divide the order of $T(\mathbb{F}_{q})$ and as $T(\mathbb{F}_{q})$ is abelian, the intersection of the kernels of all the non-trivial $k$-valued characters of $T(\mathbb{F}_{q})$ is trivial. Let $s=s_{\alpha}\in S$ be such that $|\mu_{\alpha}|=c$, and note that the group $\alpha^{\vee}(\mathbb{F}_{q}^{\times})$ is non-trivial since $q>2c$. Then there must exist a character $\xi$ with $\xi|_{\alpha^{\vee}(\mathbb{F}_{q}^{\times})}\neq 1$ and so $S_{\xi}\neq S$. ∎
@@
To see this, we use the fact that $T_{\lambda_{0}^{+}}^{n+1}=z^{n}T_{\lambda_{0}^{+}}$ (cf. [2, Lemma 5.17] and [32, Lemma 2.18]). Now let $f\in{\mathcal{R}^{G}_{M}}({\mathfrak{m}}^{\prime})$. Suppose first that $f\cdot z^{n}=0$. Then for all $h\in{\mathcal{H}}_{M}$, we have
@@
by definition of the map $j_{M}^{+}$ (cf. (2.5)). Since $T^{M}_{\lambda_{0}^{+}}$ is invertible in ${\mathcal{H}}_{M}$, this means that $f=0$. To get surjectivity, define $g:{\mathcal{H}}_{M}\to{\mathfrak{m}}^{\prime}$ by $g(h)=f(h\cdot(T^{M}_{\lambda_{0}^{+}})^{-n-1})\cdot T_{\lambda_{0}^{+}}$. This is ${\mathcal{H}}_{M^{+}}$-linear because $T^{M}_{\lambda_{0}^{+}}$ is central in ${\mathcal{H}}_{M}$, and we have
@@
${{\mathfrak{m}}}$${{\mathfrak{m}}}$${Q_{f}{\mathfrak{m}}}$${Q_{f}{\mathfrak{m}}}$$\scriptstyle{\cdot z^{n}}$$\scriptstyle{i_{{\mathfrak{m}}}}$$\scriptstyle{i_{{\mathfrak{m}}}}$$\scriptstyle{\cdot z^{n}}$
as a square in ${\operatorname{Ho}}({\mathcal{H}})$, we deduce that multiplication by $z^{n}$ induces the zero endomorphism of $Q_{f}{\mathfrak{m}}$ in the homotopy category. Thus the image of this endomorphism under the additive functor $\mathbf{R}{\mathcal{R}^{G}_{M}}$ is also zero. As $Q_{f}{\mathfrak{m}}$ is Gorenstein injective, i.e. fibrant, we have an isomorphism
$0=\mathbf{R}{\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}}\xrightarrow{\cdot z^{n}}Q_{f}{\mathfrak{m}})\cong{\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}}\xrightarrow{\cdot z^{n}}Q_{f}{\mathfrak{m}})$
in the arrow category of ${\operatorname{Ho}}({\mathcal{H}}_{M})$. As we have seen above, the right hand map is an isomorphism, from which we deduce that $\mathbf{R}{\mathcal{R}^{G}_{M}}({\mathfrak{m}})\cong{\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})\cong 0$ in ${\operatorname{Ho}}({\mathcal{H}}_{M})$, as required. ∎
@@
Similarly to the above proof, take $\lambda_{0}^{\pm}$ as in 2.1 and let $\mathcal{O}^{\pm}=\widetilde{W}\cdot\lambda_{0}^{\pm}$. Then part (i) of the above result holds more generally whenever every element of ${\mathfrak{m}}$ is killed by a power of $z_{\mathcal{O}^{-}}$, and part (ii) holds whenever $z_{\mathcal{O}^{+}}$ acts nilpotently on ${\mathfrak{m}}$. This follows from the above proof and because the abelian analogue of 3.18 holds in that higher level of generality, cf. the proof of [4, Proposition 4.27].
@@
As another corollary, we deduce that uniformly supersingular modules are Ext-orthogonal to parabolically induced modules in large enough degree without assuming any finiteness condition on these modules. This extends to any positive degree for finite length modules (cf. 3.26).
@@
Let $Q_{f}{\mathfrak{m}}$ be the fibrant replacement of ${\mathfrak{m}}$. It follows from 3.18(ii) that ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ has finite injective dimension. Since ${\mathcal{R}^{G}_{M}}(Q_{f}{\mathfrak{m}})$ is also Gorenstein injective by 3.7(iv), we deduce that it is injective by [17, Corollary 8.5]. Applying now 3.7(ii)-(iii), we obtain that
@@
Next, we compute the image of all simple modules under the adjoints of parabolic induction at the homotopy level. For that we need the corresponding formula at the level of the module categories, due to Abe, which we recall now. We refer the reader back to (2.8)-(2.9) for the construction of the simple ${\mathcal{H}}$-modules.
@@
So we may assume that $P\subseteq P^{\prime}$ and $\Pi(\mathfrak{n})\subseteq\Pi_{Q}\cup\Pi_{P^{\prime}}$. Next, for any standard parabolic $Q\cap M^{\prime}\subseteq Q^{\prime}\subseteq(P\cap M^{\prime})(\mathfrak{n})$ of $M^{\prime}$ with Levi subgroup $L^{\prime}$ we have $e_{{\mathcal{H}}_{w_{0}(L^{\prime})}}(\iota_{0}^{M}\mathfrak{n})\cong\iota_{0}^{L^{\prime}}e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n})$, cf. [4, Lemma 4.22]. From this and the above we get
@@
where the second, third and fourth isomorphisms hold by 2.3, [32, Lemma 4.9] and (2.6), respectively. By exactness of $\iota_{0}^{w_{0}(M^{\prime})}$ and from the definitions of both $I_{{\mathcal{H}}_{M^{\prime}}}(P\cap M^{\prime},\mathfrak{n},Q\cap M^{\prime})$ and $CI_{{\mathcal{H}}_{M^{\prime}}}(w_{0}(P\cap M^{\prime}),\iota_{0}^{M}\mathfrak{n},w_{0}(Q\cap M^{\prime}))$, it follows that we have an isomorphism
@@
Proposition 3.24.
Let ${\mathfrak{m}}$ be a finite length ${\mathcal{H}}$-module. Then ${\mathfrak{m}}$ is ${\mathcal{R}^{G}_{M}}$-acyclic for any standard Levi subgroup $M$ of $G$. In particular, $R^{i}{\mathcal{R}^{G}_{M}}(I_{\mathcal{H}}(P,\mathfrak{n},Q))=0$ for any $i>0$ and any standard supersingular triple $(P,\mathfrak{n},Q)$.
@@
for all $j\geq 0$, where $({\mathcal{R}^{G}_{M}})^{\prime}$ denotes the right adjoint of parabolic coinduction. By applying 3.7(iii), we deduce that finite length modules are acyclic for $({\mathcal{R}^{G}_{M}})^{\prime}$. Since $({\mathcal{R}^{G}_{M}})^{\prime}\cong\iota_{0}^{w_{0}(M)}\circ\mathcal{R}^{G}_{w_{0}(M)}$, cf. 2.3(i), and $\iota_{0}^{w_{0}(M)}$ is an exact equivalence, we get that finite length modules are acyclic for $\mathcal{R}^{G}_{w_{0}(M)}$. As $M$ was arbitrary we are done. ∎
@@
[…]
Remark 3.26.
Suppose ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ is supersingular and of finite length, and let $\mathfrak{n}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$ be any module. Then we have isomorphisms
${\operatorname{Ext}}_{{\mathcal{H}}}^{i}(I_{M}^{G}(\mathfrak{n}),{\mathfrak{m}})\cong{\operatorname{Ext}}_{{\mathcal{H}}_{M}}^{i}(\mathfrak{n},{\mathcal{R}^{G}_{M}}({\mathfrak{m}}))=0$
for all $i\geq 0$, by applying 3.24, 3.7(iii) and using the fact that ${\mathcal{R}^{G}_{M}}({\mathfrak{m}})=0$ by [2, Proposition 5.18]. This strengthens 3.21 in the finite length case.
@@
Let $M$ be a standard Levi subgroup of $G$ and let ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ be ${\mathcal{R}^{G}_{M}}$-acyclic. Then there is an isomorphism $\mathbf{R}{\mathcal{R}^{G}_{M}}({\mathfrak{m}})\cong{\mathcal{R}^{G}_{M}}({\mathfrak{m}})$ in ${\operatorname{Ho}}({\mathcal{H}}_{M})$.
@@
where $\mathfrak{e}$ has finite injective dimension. As both ${\mathfrak{m}}$ and $Q_{f}{\mathfrak{m}}$ are ${\mathcal{R}^{G}_{M}}$-acyclic, cf. 3.7(ii), so is $\mathfrak{e}$. By applying 3.7(iii), we then get isomorphisms
@@
The statement for the left adjoints follows immediately from 3.23. The statement for the right adjoints follows additionally from 3.24 and 3.27. ∎
It was shown in [21, Theorem 8.5] that, when $p\nmid\left|\Omega_{\text{tor}}\right|$, a simple module $I_{\mathcal{H}}(P,\mathfrak{n},Q)$ has finite projective dimension if and only if $\mathfrak{n}$ has finite projective dimension. We now give a new proof of the ‘only if’ direction without any assumptions on the prime $p$ and without using any Bruhat-Tits theory. This result is not new, in that the proof of the equivalence in loc. cit. only uses their assumption on $p$ in the ‘if’ direction, although they never state this explicitly. Using their restriction on $p$, we can also essentially reduce our aim of studying the isomorphism classes of simple modules in ${\operatorname{Ho}}({\mathcal{H}})$ to the supersingular case.
@@
    Assume that $p\nmid\left|\Omega_{\text{tor}}\right|$ and that ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})={\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}}^{\prime})=\infty$. If ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ then $P=P^{\prime}$, $Q=Q^{\prime}$ and $\mathfrak{n}\cong\mathfrak{n}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}}_{M})$.
@@
(i) Let $L$ denote the Levi subgroup of $Q$. Then we have $0=\mathbf{R}\mathcal{R}^{G}_{L}(I_{\mathcal{H}}(P,\mathfrak{n},Q))\cong I_{{\mathcal{H}}_{L}}(P\cap L,\mathfrak{n},L)$ by 3.28, so that we may assume that $Q=G$. Applying the left adjoint instead, we now have
@@
(ii) First suppose that $Q\neq Q^{\prime}$. Up to swapping them we may assume that $Q\nsubseteq Q^{\prime}$. Write $L^{\prime}$ for the Levi subgroup of $Q^{\prime}$. Then applying $\mathbf{R}\mathcal{R}^{G}_{L^{\prime}}$ to ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$, we obtain $0\cong I_{{\mathcal{H}}_{L^{\prime}}}(P^{\prime}\cap L^{\prime},\mathfrak{n}^{\prime},L^{\prime})$ by 3.28, contradicting our assumption that ${\mathfrak{m}}^{\prime}$ has infinite projective dimension by [21, Theorem 8.5]. Thus $Q=Q^{\prime}$, and applying $\mathbf{R}\mathcal{R}^{G}_{L^{\prime}}$ again we may assume that $Q=Q^{\prime}=G$ as above. Similarly now, if $P\neq P^{\prime}$ then without loss of generality we have $P\nsubseteq P^{\prime}$ and, writing $M^{\prime}$ for the Levi subgroup of $P^{\prime}$, we then obtain a contradiction from 3.28 by applying ${\operatorname{Ho}}(\mathcal{L}^{G}_{M^{\prime}})$. ∎
@@
Suppose that $p\nmid\left|\Omega_{\text{tor}}\right|$, and let $\mathfrak{s}=I_{\mathcal{H}}(P,\mathfrak{n},Q)\in{\operatorname{Mod}}({\mathcal{H}})$ be a simple module of infinite projective dimension. Then we have:
@@
First observe that for any standard parabolic subgroup $P^{\prime}=M^{\prime}N^{\prime}$ of $G$, we have that $I_{{\mathcal{H}}_{M^{\prime}}}(P,\mathfrak{n},Q\cap P^{\prime})$ and $I_{{\mathcal{H}}_{M^{\prime}}}(P,\mathfrak{n},Q)$ also have infinite projective dimension, whenever these modules are well-defined. Indeed, this follows from [21, Theorem 8.5] by our assumption on $p$ and because $\mathfrak{s}$ is assumed to have infinite projective dimension. The statement of (ii) then follows from 3.28 and 3.13(iii), and for (i) we get
@@
Suppose that $p\nmid\left|\Omega_{\text{tor}}\right|$. When the root system of $G$ is irreducible, the following are equivalent for an ${\mathcal{H}}$-module ${\mathfrak{m}}$ of finite length:
@@
    the inclusion ${\mathfrak{m}}_{\text{ss}}\subseteq{\mathfrak{m}}$ is a weak equivalence, where ${\mathfrak{m}}_{\text{ss}}$ denotes the largest supersingular submodule of ${\mathfrak{m}}$.
@@
Suppose first that $\mathfrak{s}=I_{\mathcal{H}}(P,\mathfrak{n},Q)$ is a simple subquotient of ${\mathfrak{m}}_{\text{non-ss}}$. If $P(\mathfrak{n})\neq G$ then $\mathfrak{s}$ is in the essential image of $I^{G}_{M(\mathfrak{n})}$. If instead $P(\mathfrak{n})=G$ then $P=B$ since $P\neq G$ and the root system is irreducible, and thus $\mathfrak{s}$ has finite projective dimension by [21, Theorem 8.5] and the fact that ${\mathcal{H}}_{T}$ has finite global dimension (cf. 3.6). Either way, it follows that $\mathfrak{s}\in{\mathcal{S}}^{G}_{\text{all}}$ and hence that ${\mathfrak{m}}_{\text{non-ss}}\in{\mathcal{S}}^{G}_{\text{all}}$.
@@
When the hypotheses of 3.32 are satisfied we tentatively call the quotient category ${\operatorname{Ho}}({\mathcal{H}})_{\text{all}}$ the supersingular homotopy category of ${\mathcal{H}}$. We will see later that this triangulated category is in fact the homotopy category of a model structure on ${\operatorname{Mod}}({\mathcal{H}})$.
@@
If we write ${\mathcal{T}}$ for ${\operatorname{Ker}}(\mathcal{L}^{G}_{\text{all}})$, then 3.32 tells us that the modules of finite length in ${\mathcal{T}}$ are, up to weak equivalence, the supersingular ones whenever the root system of $G$ is irreducible and $p\nmid\left|\Omega_{\text{tor}}\right|$. Moreover we have ${\mathcal{T}}^{\ddagger}={\mathcal{S}}^{G}_{\text{all}}$ and ${\mathcal{T}}={{}^{\ddagger}}{\mathcal{S}}^{G}_{\text{all}}$. This is similar to the torsion pair constructed at the abelian level in [28] where the authors showed that when $\mathbb{G}=\text{SL}_{2}$, $p>2$ and $\mathfrak{F}\neq\mathbb{Q}_{p}$, then the supersingular modules are in a torsion pair with a certain class of modules which they call ‘non-supersingular’ and which contains all parabolically induced modules, cf. Theorems 3.26 & 3.27 in loc. cit. Since our Hom-orthogonality boils down to Ext-orthogonality between suitable replacements at the abelian level, the two questions are not immediately related. But we hope to study elsewhere the relationship between the Ollivier-Schneider torsion pair and the adjoints of parabolic induction.